OpenStudy (anonymous):

Use Stokes' Theorem to evaluate the surface integral where F= and where S is the portion of z=1-x^2-y^2 above the xy-plane with n upward. (Answer: -pi)

OpenStudy (anonymous):

Haven't started Calc III yet =/ You are done Calc II right? @Idealist

OpenStudy (anonymous):

I hate these questions because it's ambiguous what they mean by \(\mathbf F\)

OpenStudy (anonymous):

\[ \iint_S (\nabla \times \mathbf F) \cdot d\mathbf S = \int_{\partial S}\mathbf F\cdot d\mathbf s \]

OpenStudy (anonymous):

See the ambiguity is whether our vector is post-curl or pre-curl. In this case, I'm assuming \(\mathbf F\) is going to be pre-curl.

OpenStudy (anonymous):

@Idealist Can you draw the surface?

OpenStudy (anonymous):

This looks like a circle. x^2+y^2=1, where r=1. So x=cos t, y=sin t, z=0 and dx=-sin t, dy=cos t, dz=0. And take the integral of zx^2 dx+z(e^xy^2)-x dy+x*ln(y^2) dz and I got 0 as the answer, which doesn't match the answer in the book. So can someone please correct me?

OpenStudy (anonymous):

Don't they want you to take the curl instead?

OpenStudy (anonymous):

The curl is <2y/y^2-1, 0, ze^y^2-1> but what do I do with that?

OpenStudy (anonymous):

You can integrate over any surface with the same boundary.

OpenStudy (anonymous):

that is another kicker of stokes theorem.

OpenStudy (anonymous):

But how do I set up the integral with the limits?

OpenStudy (anonymous):

Is this parametrisation correct though? I do believe that all points corresponding in the relation to the radius are important, so shouldn't it be: x=rcost y=rsint z=1-r^2 while r ranges from 0 to 1 and t ranges from 0 to 2 pi

OpenStudy (anonymous):

@Spacelimbus He parametrized the boundary, not the actual surface.

OpenStudy (anonymous):

I need to look up on that then, seems complicated though, not suggesting that it's wrong then.

OpenStudy (anonymous):

But should I find the dS?

OpenStudy (anonymous):

@Idealist What did you get for \(\mathbf F\circ \mathbf r(t)\)?

OpenStudy (anonymous):

I wanna see a step at a time.

OpenStudy (anonymous):

What's r(t)?

OpenStudy (anonymous):

Your parametrization of the boundary in this case.

OpenStudy (anonymous):

I really don't know how to find that.

OpenStudy (anonymous):

You have to find it to do a line integral... composing the function with the parametrization... You gotta know how to do that.

OpenStudy (anonymous):

\[ \mathbf F(x(t),y(t),z(t)) \]

OpenStudy (anonymous):

Oh, and you find dx, dy, dz with the given, right?

OpenStudy (anonymous):

Yes. I want to see your work.

OpenStudy (anonymous):

Wait a minute.

OpenStudy (anonymous):

dx=2xz, dy=ze^2yx, dz=0.

OpenStudy (anonymous):

Nope, \(dx = x'(t)dt\)

OpenStudy (anonymous):

Just compose this function with the parametrizaiton first... can I see that?\[ \mathbf F=<zx^2, z(e^xy^2)-x, x*ln(y^2)> \]

OpenStudy (anonymous):

No, it's z(e^(xy^2)-x),

OpenStudy (anonymous):

\[ \mathbf F=<zx^2, z(e^{xy^2}-x), x\ln(y^2)> \]

OpenStudy (anonymous):

dx=2xz dt dy=ze^(2yx) dt dz=0 dt

OpenStudy (anonymous):

Wrong.

OpenStudy (anonymous):

Are you finding the gradient? Why?

OpenStudy (anonymous):

Sorry, I'm really confused by what you are trying to do.

OpenStudy (anonymous):

As I said before, I hate these problems, because they somewhat specific to the book.

OpenStudy (anonymous):

@Idealist Can you just ignore the whole \(dx,dy,dz\) part and just compose the vector field with the parametrization first? Please?

OpenStudy (anonymous):

Yes, I'll try. I think I need to look up online to see the procedure because my book might be wrong, sorry to make you confused. And thank you for the help.

OpenStudy (anonymous):

To do the composition, you have to replace the \(x,y,z\) with their respective functions of \(t\). There should be no \(x,y,z\).